ADNAN MUHAMMED

Grade 12,

An electron at rest has a charge of 1.6 × 10_19 C. It starts moving with a velocity v = c/2, where c is the speed of light, then the new charge on it is

An electron at rest has a charge of 1.6 × 10_19 C. It starts moving with a velocity v = c/2, where c is the speed of light, then the new charge on it is

Grade:12

2 Answers

Vikas TU
6 years ago
Amount of charge = 1.6 x 10^-19 CVelocity = c/2        = 1.5 x 10^6 m/sec Then the new charge = 1.6 x 10^-19 x 1.5 x 10^6                     = 2.4 x 10^-13 C
ankit singh
3 years ago
Answer:2.4*10^-13 C
Explanation:Dear Student
According to the question:The amount of charge = 1.6*10^(-19)C
Velocity  =c/2= 1.5*10^6 m/sec.
Then the new charge = 1.6*10^-19*1.5*10^6 C
                                     = 2.4*10^-13 C